OpenStudy (anonymous):

What is the distance between (6, 2) and (-3, -2)?

OpenStudy (gorv):

use the diatance formula

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 6}}\quad ,&{\color{blue}{ 2}})\quad &({\color{red}{ -3}}\quad ,&{\color{blue}{ -2}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (gorv):

\[\sqrt{(x2--x1)^{2}+(y2-y1)^{2}}\]